What are the focus and the directrix of the graph of x = 1/24y^2?
focus (0, –6), directrix x = 6
focus (6, 0), directrix x = –6
focus (0, 6), directrix y = –6
focus (–6, 0), directrix y = 6

Answer by lwsshak3(11628)

What are the focus and the directrix of the graph of x = 1/24y^2?
y^2/24=x
y^2=24x
This is an equation of a parabola that opens rightward:
Its basic equation: (y-k)^2=4p(x-h), (h,k)=(x,y) coordinates of the vertex
For given equation:
vertex:(0,0)
axis of symmetry: x-axis
4p=24
p=6
focus: (6,0) (p-distance to the right of the vertex on the axis of symmetry)
directrix: x=-6 (p-distance to the left of the vertex on the axis of symmetry)